# SwiftKV: Fast Prefill-Optimized Inference with Knowledge-Preserving Model Transformation

### v2

@@ Appendix C Inter-layer AcrossKV vs Intra-Layer KV cache Reduction
In this section, we share different design choices of AcrossKV, which considers the tradeoff between GQA (Ainslie et al., 2023a) and the acorss layer sharing into the design. Particularly, when $\text{AcrossKV}\geq 2$, we can either use GQA and AcrossKV together or we can simply use AcrossKV to get all savings. For instance, when $\text{AcrossKV}=4$ (a.k.a., the second row of the final session in Table 2), we have KV cache reduction from both GQA and AcrossKV. However, we can either do multi-query attention (MQA) for all 16 layers or do multi-head attention (MHA) but share the KV cache for all 16 layers.
Table C.1: Llama-3.1-8B-Instruct AcrossKV design
@@
We present the $50\%$ SingleInputKV reduction with MQA, GQA plus AcrossKV, and GQA plus MHA in Table C.1, that all have the same KV cache reduction, 37.5%. AcrossKV-GQA actually provides the best performance. One thing to notice is that the AcrossKV-MHA is actually worse than the result of AcrossKV-16x from from Table 2 even though AcrossKV-MHA has larger KV cache than AcrossKV-16x. We hypothesis that this might be related to hyper-paramter tuning but did not invest deeper. Also, note that pure MQA leads to worst performance, which is about 17 points lower than AcrossKV-GQA
@@
Appendix D The impact of fine-tuning datasets
@@
How good is the data used to train SwiftKV? We chose the datasets to train SwiftKV due to their popular adoption and broad domain and task coverage. However, as compared to other high-quality domain specific fine-tuning datasets, they may have weaknesses. To measure the quality of these two datasets, we directly fine-tuned a model using the Llama-3.1-8B base model, and compared this trained model with the Llama-3.1-8B-Instruct model released by Meta.
The results are shown in Table D.1 (a). The original Llama-3.1-8B-Instruct has a average score of 73.71 but the model trained using our two datasets only achieved 65.77. This indicates the training data used for SwiftKV is not optimal and there may be opportunities to further improve the results we reported in Sec. 4 as discussed next.
Does more math/coding data help GSM-8K? From Table 2, the main degradation among 7 tasks for 50% SingleInputKV is GSM-8K. This may be due to the lack of math and coding examples in the two datasets we picked to train the model. To verify this, we distilled SwiftKV using one extra math-related dataset, gretelai/synthetic-gsm8k-reflection-405b (GretelAI, 2024), and one extra coding dataset, ise-uiuc/Magicoder-OSS-Instruct-75K (Wei et al., 2023), in total about $8\text{K}+75\text{K}=83\text{K}$ samples, and about 16M tokens.
The results are reported in Table D.1 (b). The performance of all tasks except Winogrande are slightly improved, with the average score being 0.23 higher. Particularly, GSM-8K improves the most, with a 0.53% improvement. This is expected since we added extra math and coding datasets. Considering the small amount of new data (83k vs. 1.2M), the improvement is remarkable.
@@
Table D.1: The impact of datasets on Llama-3.1-8B-Instruct.
@@
Figure D.1: Density of early exit probabilities and alignment of early exit vs final logits.
@@ 3. Yellow
Table D.2: A Q&A example of early exit.
Appendix E Simple Early Exit for Decoding Tokens
SingleInputKV allows all the KV cache needed for generating future tokens to be computed without having to forward-pass though the entire LLM. This means that even the decoding phase could exit earlier without worrying about missing KV cache for subsequent tokens.
To test the feasibility, we added an early exit language modeling head. We then used the input to SingleInputKV layer to calculate output logits, and incorporated them as part of the distillation training. Our results are preliminary and requires further evaluation, but we found that the alignment between early exit logits and the final output logits to be over 66% when the largest probability from the early exit logits is over 95% (Fig. D.1). We used this as our early exit criteria (i.e., decoding tokens exit early if it predicts an output token with 95%+ probability), and  Table D.2 shows a sample result. See Appendix F for more details.
Appendix F Early Exit Details
Thanks to SingleInputKV, there is no need to go through the entire network to compute all KV cache for later tokens generation. This also brings another unique opportunity as compared to standard transformer architecture design: it allows the model to perform early exit to speedup both the prefill and generation phases without worrying about missing KV cache.
To add early exit upon SwiftKV, we add an extra language modeling head and use the input to calculate SingleInputKV to directly compute the output logits. We apply the same distillation loss of this early exit logits with the original teacher’s logits and directly sum the two distillation losses together as our final optimizing loss.
After training, we first look the alignment between the early exit logits and the final logits. The figure is shown in Fig. D.1. The left axis plot the density of early exit largest probabilities. And the right axis shows the probabilities when the maximum of early exit logits aligns with the final logits for each bin. As can be seen, when the maximum logits is larger than $0.95$, the alignment likelihood is larger than 66%. Note that this is achieved within 160M token training and longer training should help the alignment.
How to use early exit is always an interesting directly and research topic. Fully exploring this is out of the scope of this paper. Here we adopt a simple heuristic approach. When the largest probability is larger than 0.95, we directly use the early exit prediction. Otherwise, we still process the remaining layers to get the final prediction. We apply this logic to some questions we selected from Alpaca (Taori et al., 2023) and the results are shown in Table D.2 and Appendix F.1.
F.1 Early Exit Examples

### v3

@@ B.2 Inter-layer AcrossKV vs Intra-Layer KV cache Reduction
In this section, we share different design choices of AcrossKV, which considers the tradeoff between GQA (Ainslie et al., 2023a) and the across layer sharing into the design. Particularly, when $\text{AcrossKV}\geq 2$, we can either use GQA and AcrossKV together or we can simply use AcrossKV to get all savings. For instance, when using 4$\times$ AcrossKV, we have KV cache reduction from both GQA and AcrossKV. However, we can either do multi-query attention (MQA) for all 16 layers or do multi-head attention (MHA) but share the KV cache for all 16 layers.
Table B.2: Llama-3.1-8B-Instruct AcrossKV design
@@
We present the $50\%$ SwiftKV reduction with MQA, GQA plus AcrossKV, and GQA plus MHA in Table B.2, that all have the same KV cache reduction, 37.5%. AcrossKV-GQA actually provides the best performance. One thing to notice is that the AcrossKV-MHA is actually worse than the result of 16$\times$ AcrossKV from from Table 2 even though AcrossKV-MHA has larger KV cache than 16$\times$ AcrossKV. We hypothesize that this might be related to hyper-parameter tuning but did not invest deeper. Also, note that pure MQA leads to worst performance, which is about 17 points lower than AcrossKV-GQA
@@
B.3 The impact of fine-tuning datasets
@@
How good is the data used to train SwiftKV?
We chose the datasets to train SwiftKV due to their popular adoption and broad domain and task coverage. However, as compared to other high-quality domain specific fine-tuning datasets, they may have weaknesses. To measure the quality of these two datasets, we directly fine-tuned a model using the Llama-3.1-8B base model, and compared this trained model with the Llama-3.1-8B-Instruct model released by Meta.
The results are shown in Table B.3 (a). The original Llama-3.1-8B-Instruct has a average score of 73.71 but the model trained using our two datasets only achieved 65.77. This indicates the training data used for SwiftKV is not optimal and there may be opportunities to further improve the results we reported in Sec. 4 as discussed next.
Does more math/coding data help GSM-8K?
From Table 2, the main degradation among 7 tasks for 50% SwiftKV is GSM-8K. This may be due to the lack of math and coding examples in the two datasets we picked to train the model. To verify this, we distilled SwiftKV using one extra math-related dataset, gretelai/synthetic-gsm8k-reflection-405b (GretelAI, 2024), and one extra coding dataset, ise-uiuc/Magicoder-OSS-Instruct-75K (Wei et al., 2023), in total about $8\text{K}+75\text{K}=83\text{K}$ samples, and about 16M tokens.
The results are reported in Table B.3 (b). The performance of all tasks except Winogrande are slightly improved, with the average score being 0.23 higher. Particularly, GSM-8K improves the most, with a 0.53% improvement. This is expected since we added extra math and coding datasets. Considering the small amount of new data (83k vs. 1.2M), the improvement is remarkable.
@@
Table B.3: The impact of datasets on Llama-3.1-8B-Instruct.
@@
Figure B.1: Density of early exit probabilities and alignment of early exit vs final logits.
@@ 3. Yellow
Table B.4: A Q&A example of early exit.
B.4 Simple Early Exit for Decoding Tokens
SwiftKV allows all the KV cache needed for generating future tokens to be computed without having to forward-pass though the entire LLM. This means that even the decoding phase could exit earlier without worrying about missing KV cache for subsequent tokens.
To add early exit upon SwiftKV, we add an extra language modeling head and use the input to calculate SwiftKV to directly compute the output logits. We apply the same distillation loss of this early exit logits with the original teacher’s logits and directly sum the two distillation losses together as our final optimizing loss.
After training, we first look the alignment between the early exit logits and the final logits. The figure is shown in Fig. B.1. The left axis plot the density of early exit largest probabilities. And the right axis shows the probabilities when the maximum of early exit logits aligns with the final logits for each bin. As can be seen, when the maximum logits is larger than $0.95$, the alignment likelihood is larger than 66%. Note that this is achieved within 160M token training and longer training should help the alignment.
How to use early exit is always an interesting directly and research topic. Fully exploring this is out of the scope of this paper. Here we adopt a simple heuristic approach. When the largest probability is larger than 0.95, we directly use the early exit prediction. Otherwise, we still process the remaining layers to get the final prediction. We apply this logic to some questions we selected from Alpaca (Taori et al., 2023) and the results are shown in Table B.4 and Appendix B.4.1.
B.4.1 Early Exit Examples